OpenStudy (anonymous):

8x^2-2x-3=0, please tell me the way how you did it.

OpenStudy (anonymous):

believe it or not this one factors, but you can always use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] if it is not clear how to factor it

OpenStudy (anonymous):

Factor to get: \[(4x-3)(2x + 1)\ =\ 0\]

OpenStudy (anonymous):

\[4x -3 = 0\ \ OR\ \ 2x +1 =0\] \[x = \frac{3}{4}\ \ or \ x = -\frac{1}{2} \]

OpenStudy (anonymous):

This one could be done by grouping i believe. 8 times -3 is -24, and the middle number is -2, so you want to think of two numbers that multiply to -24 and add to -2. These numbers are -6 and 4. So now we change that -2x to (-6x+4x) and we get: \[8x^2-2x-3 = 8x^2+4x-6x-3 = 4x(2x+1)-3(2x+1) \] \[= (4x-3)(2x+1)\]

OpenStudy (anonymous):

i love "factor by grouping" but i can never to it!

OpenStudy (mathteacher1729):

You could try factoring: Here is my thought process: (8x + a ) ( x + b ) = 8x^2-2x-3 or (2x + a)( 4x + b) = 8x^2-2x-3 Now the choices of a and b can be anything in the set { +1 , -1 , +3 , -3}

OpenStudy (anonymous):

i learned how to do it on this site lol, someone else said what a posted and i was like "OH SO THATS HOW YOU DO IT!!" lolol

OpenStudy (anonymous):

hey joemath you are completely right but how did u get 4x,-6x?

OpenStudy (anonymous):

i got it men thank u.